Seafoam:

Ellis is painting wooden fenceposts before putting them in his yard. They are each 5 feet tall and have a diameter of 1 foot. There are 10 fenceposts in all. How much paint will Ellis need to paint all the surfaces of the 10 fenceposts? Use 3.14 for π, and round your answer to the nearest hundredth. Provide an explanation and proof for your answer.

Seafoam:

@AZ

AZ:

so they're cylinders and we're finding the surface area for one and multiplying by 10 since there's 10 fenceposts

Seafoam:

So 39.3?

AZ:

SA = (2* pi * radius * height) + 2*pi*(radius)^2 diameter = 1 ft radius = diameter / 2 so what is the radius? and height = 5 and use 3.14 for pi

Seafoam:

I thought .5 was the radius?

AZ:

It is!

Seafoam:

In which case I already did that all and I got 39.3

Seafoam:

This is so much easier when explained

AZ:

hmm that's after you're multiplying by 10, I guess? I'm getting a slightly different answer

Seafoam:

Ok I put .5 as the radius and 5 as the height, it said the volume was 3.93 and mulitplied by 10 thats 39.3

AZ:

wait I'm dumb

AZ:

I was accidentally using r= 0.1 instead of 0.5

AZ:

But I don't get the same answer as you though Let's plug in the numbers, radius = 0.5 and height = 5 SA = (2* pi * radius * height) + 2*pi*(radius)^2 SA = (2*3.14*0.5*5) + (2*3.14*0.5*0.5) SA = ??

Seafoam:

Oh

Seafoam:

You're solving for surface area...

Seafoam:

The default setting is volume

Seafoam:

whoops

Seafoam:

And now it says 174.36

AZ:

that makes sense haha

AZ:

oops be careful that you're doing 0.5 * 0.5

Seafoam:

So just times ten

AZ:

and use 3.14 for pi

Seafoam:

Oh wait what?

AZ:

SA = (2*3.14*0.5*5) + (2*3.14*0.5*0.5)

Seafoam:

Ok I kind of just searched up volume of a cylinder to help myself out

AZ:

surface area*** just use a calculator and plug in the numbers for the first parenthesis and then the second one and add you were on the right track but just forgot the decimal in the radius- 0.5 and not 5

Seafoam:

Oh my god

Seafoam:

I put 55 instead of 5

Seafoam:

*head slap*

AZ:

ye, I accidentally made a typo 3 times too when calculating it so no worries

Seafoam:

And so 172.8

Seafoam:

Thats with it multipled by 10

AZ:

Yup! I got 172.7 but you were probably using the actual value of pi instead of just 3.14

Seafoam:

So its rounded up to 172.9

AZ:

no no just 172.7 is fine

AZ:

and the units would be square feet

Seafoam:

A=2πrh+2πr2 R= .5 H= 5 A=2πrh+2πr2=2·π·0.5·5+2·π·0.52≈17.27876 SA= 17.28 17.28 times 10 =  172.8

Seafoam:

Does that work?

AZ:

yeah, except that they want you to use 3.14 for pi `Use 3.14 for π` so just say that the final answer is 17.27 and 17.27 * 10 = 172.70 added an extra zero because `round your answer to the nearest hundredth`

Seafoam:

A=2πrh+2πr2 R= .5 H= 5 A=2πrh+2πr2=2·π·0.5·5+2·π·0.52≈17.27876 SA= 17.28 17.27 * 10 = 172.70

AZ:

yeah I guess that works haha

Seafoam:

yay ok, next